OpenStudy (anonymous):

What is the ratio of the area of the pool to the area of the deck? Pool: 20*60=1200ft^2 Deck: 30*90=2700ft^2

OpenStudy (anonymous):

A ratio is a fraction. Pool/Deck So, 1200ft^2/2700ft^2 1200/2700 = 12/27

OpenStudy (anonymous):

Or 44.44 %

OpenStudy (anonymous):

okay. So how many feet of fencing is needed to go around the deck?

OpenStudy (anonymous):

Fencing problems refer to perimeter. To calculate perimeter you simply add all four sides together. If the deck is 30*90, then there are 2 sides of length 30 and 2 sides of length 90. Therefor, 2(30) + 2(90) = 240ft